OpenStudy (anonymous):

the perimeter of a sign must be 114 inches. Also its length is 13 inches longer than its width. Find the dimensions of this sign. The width of the sign is? The length of the sign is?

OpenStudy (anonymous):

p=2l+2w 114=2(w+13)+2w 57=w+13+w 57-13=2w 44=2w w=22 22*2=44 +35*2=114 Therefore width of 22 and length of 35